# Ultracold Interactions between Ions and Polar Molecules

### v1

@@ 7.1 Control of collisional processes in molecular gases
Applying sufficiently large homogeneous electric fields along the strongly confining axis has been shown to suppress short-range collisions by preventing head-to-tail collisions [81]. In the proposed ion-molecules experiments a similar arrangement can be realized by replacing the external electric field with a single ion positioned between two lattice sites, as illustrated in Fig. 9. Direct contact between the ion and the molecules would be avoided as long as the ion’s position can be controlled on a sub-wavelength scale. This capability was demonstrated in several early works [101, 102]. For a common wavelength of a NIR dipole trap of $\lambda=1064~{}\textrm{nm}$, the electric field of an ion displaced by $\lambda/4$, that is, 0.5 of the lowest lattice site spacing, would lead to an equally small induced dipole moment of NaK in two adjacent sites of $0.03~{}d_{0}$. Considering the van der Waals potential between two molecules, the effective barrier stemming from induced dipole-dipole repulsion is then $U_{0}/k_{B}=234~{}\textrm{nK}\sim T_{mol}$. This is on the order of the temperature in typical non-degenerate molecular quantum gases such that no significant change of the inelastic collision rate is expected. However, moving the ion closer to a lattice site gives rise to a substantial barrier. As an example, at an ion-lattice site spacing of $d_{lat}=(\lambda/4,~{}\lambda/8,~{}\lambda/16)\approx(133,~{}66,~{}33)~{}\textrm{nm}$ the barrier height increases to $U_{0}/k_{B}=(0.06,~{}6.2,~{}52.1)~{}\textrm{mK}\gg T_{mol}$. This suggests that ion-mediated shielding of short-range collisions may also be achieved using an ion as a nanoscale source of external electric fields without immersing it into the molecular cloud as discussed in section 6. An additional advantage of this approach is that changing $d_{lat}$, simultaneously lowers the barrier in the adjacent lattice site, e.g. to $U_{0}/k_{B}<10~{}\textrm{nK}\ll T_{mol}$ for $d_{lat}=\lambda/4$ to $\lambda/8$. In this fashion systematic effects between different realizations of an experiment can be reduced allowing for precise differential measurements of the influence of electric fields on the dynamical properties of molecular ensembles such as rates elastic and reactive or sticky collisions.
Figure 9: Ion as a high precision source of electric fields. Ba+ is positioned between the sites of a 1D lattice of NaK ensembles using a Paul trap or an optical dipole trap. The ion-lattice site distance $d_{lat}$ can be controlled with sub-wavelength precision [102].
7.2 Application to quantum technologies
@@
In general, a detailed understanding of solubility processes on the quantum level is of high practical importance, e.g. for the development of drugs. One way to harness the advantage of quantum computation already with the limited computational resources of platforms available today, is to map specific problems onto simplified models that can be implemented on digital quantum simulators [103]. In this context, ions immersed into molecular quantum gases closely resemble atomic cations embedded in a dense polarizable medium, a commonly encountered system. With the expected degree of control this hybrid platform may allow highly precise analog quantum simulations of more complex solubility processes that are currently studied using digital quantum simulators [104].
@@
From the perspective of quantum information processing, atomic ions are among the most performative and advanced platforms in the field [85]. Complementary to these developments, the tremendous progress achieved in the control over ultracold molecules has also brought forward promising protocols for realizing gate operations by exploiting dipole-dipole interactions [105], or by making use of the internal rotational states and their coupling to microwave fields to implement and manipulate qudits [106]. More recent ideas involve atomic Rydberg atoms to establish and distribute entanglement between molecules [107, 108]. In addition to these prominent examples, hybrid approaches promise to combine the most advantageous properties of different experimental systems. Along these lines, the combination of ions and polar molecules may prove a promising extension to the quantum toolbox of platforms using ions, atoms or molecules. As an example, in comparison to quantum computation protocols based on the interaction between ions and neutral atoms [18], a hybrid approach motivated by early experiments on ion-atom collisions, the long-range interaction between an ion and polar molecules is enhanced by the ratio of their respective static polarizabilities, that is by a factor of typically $10^{2}$ to $10^{4}$. For instance, at field strength above $\mathcal{E}^{\ast}$ corresponding to ion-neutral distances $>100~{}\mathrm{nm}$ the long-range interaction between Ba+ and NaK is about 3000 times larger than the charge-induced dipole interaction with a Rb atom. These properties may be used to realize ion-molecule interfaces or quantum computation schemes exploiting the dependence of forces exerted on a molecule by an ion on the rotational state of the molecule or the electronic state of the ion, e.g. the S ground state or a Rydberg state.
8 Conclusion
@@
In summary, we have discussed the potential realization of a hybrid ion-molecule system suited for studies of their interactions in the quantum dominated regime. We expect that the combination of currently available techniques and methodology from the fields of ion-atom interactions, in particular bichromatic dipole traps and optimized rf-based hybrid apparatus as well as protocols for the generation of ground-state polar molecules, is feasible within the scope of the envisioned studies. We identify and discuss several promising research directions in different regimes of the molecular ensembles, trap configurations and geometries. Such investigations include reactive inter-species processes with several open channels for chemical reactions and the creation of polyatomic molecular ions. Based on our calculations, we expect that despite the seemingly much higher complexity compared to ion-atom or neutral molecule systems, the additional degrees of freedom, foremost the rotational states of the molecules, can be exploited and controlled to an extent allowing for the realization of tailored schemes for ion-mediated shielding of reactive molecule-molecule and molecule-ion collisions, as well as sympathetic cooling. Furthermore, we found indications that the interplay between attractive and repulsive interactions induced by the presence of the ion in a dense polar medium facilitates the formation of many-body self-bound states.
@@
These capabilities open the way to access the elusive regime of quantum interactions between ions and neutrals holding great promise to advance several fields and applications including refined control of collisional properties in ensembles of polar molecules, quantum chemistry, quantum-many-body physics, as well as quantum simulations and technologies.

### v2

@@ VI.1 Control of collisional processes in molecular gases
Applying sufficiently large homogeneous electric fields along the strongly confining axis has been shown to suppress short-range collisions by preventing head-to-tail collisions [98]. In the proposed ion-molecules experiments a similar arrangement can be realized by replacing the external electric field with a single ion positioned between two lattice sites. The capability to control the ion-lattice site distance $d_{lat}$ with the sub-wavelength precision necessary for avoiding direct contact between the ion and the molecules was demonstrated in several early works [118, 119]. For a common wavelength of a NIR dipole trap of $\lambda=1064~{}\textrm{nm}$, the electric field of an ion displaced by $\lambda/4$, that is, 0.5 of the lowest lattice site spacing, would lead to an equally small induced dipole moment of NaK in two adjacent sites of $0.03~{}d_{0}$. Considering the van der Waals potential between two molecules, the effective barrier stemming from induced dipole-dipole repulsion is then $U_{0}/k_{B}=234~{}\textrm{nK}\sim T_{mol}$. This is on the order of the temperature in typical non-degenerate molecular quantum gases such that no significant change of the inelastic collision rate is expected. However, moving the ion closer to a lattice site gives rise to a substantial barrier. As an example, at an ion-lattice site spacing of $d_{lat}=(\lambda/4,~{}\lambda/8,~{}\lambda/16)\approx(133,~{}66,~{}33)~{}\textrm{nm}$ the barrier height increases to $U_{0}/k_{B}=(0.06,~{}6.2,~{}52.1)~{}\textrm{mK}\gg T_{mol}$. This suggests that ion-mediated shielding of short-range collisions may also be achieved using an ion as a nanoscale source of external electric fields without immersing it into the molecular cloud as discussed in Section V.2. An additional advantage of this approach is that changing $d_{lat}$ simultaneously lowers the barrier in the adjacent lattice site, e.g. to $U_{0}/k_{B}<10~{}\textrm{nK}\ll T_{mol}$ for $d_{lat}=\lambda/4$ to $\lambda/8$. In this fashion, systematic effects between different realizations of an experiment can be reduced allowing for precise differential measurements of the influence of electric fields on the dynamical properties of molecular ensembles such as the rates for elastic and reactive or sticky collisions.
VI.2 Application to quantum technologies
@@
One way to harness the advantage of quantum computation already with the limited computational resources of platforms available today, is to map specific problems onto simplified models that can be implemented on digital quantum simulators [120]. For instance, a detailed understanding of solubility processes on the quantum level is of high practical importance, e.g. for the development of drugs. In this context, ions immersed into molecular quantum gases closely resemble atomic cations embedded in a dense polarizable medium, a commonly encountered system. With the expected degree of control this hybrid platform may allow highly precise analog quantum simulations of more complex solubility processes that are currently studied using digital quantum simulators [121].
@@
From the perspective of quantum information processing, atomic ions are among the most advanced high-performance platforms in the field [105]. Complementary to these developments, the tremendous progress achieved in the control over ultracold molecules has also brought forward promising protocols for realizing gate operations by exploiting dipole-dipole interactions [122], or by making use of the internal rotational states and their coupling to microwave fields to implement and manipulate qudits [123]. More recent ideas involve Rydberg atoms to establish and distribute entanglement between molecules [124, 125]. In addition to these prominent examples, hybrid approaches promise to combine the most advantageous properties of different experimental systems. Along these lines, the combination of ions and polar molecules may prove a promising extension to the quantum toolbox of platforms using ions, atoms or molecules. As an example, in comparison to quantum computation protocols based on the interaction between ions and neutral atoms [20], a hybrid approach motivated by early experiments on ion-atom collisions, the long-range interaction between an ion and polar molecules is enhanced by the ratio of their respective static polarizabilities, that is by a factor of typically $10^{2}$ to $10^{4}$. For instance, at field strength above $\mathcal{E}^{\star}$ corresponding to ion-neutral distances $<100~{}\mathrm{nm}$ the long-range interaction between Ba+ and NaK is about 3000 times larger than the charge-induced dipole interaction with a Rb atom. These properties may be used to realize ion-molecule interfaces or quantum computation schemes exploiting the dependence of forces exerted on a molecule by an ion on the rotational state of the molecule or the electronic state of the ion, e.g. the S ground state or a Rydberg state.
VII Conclusion
@@
In summary, we have discussed the potential realization of a hybrid ion-molecule system suitable for studies of their interactions in the quantum dominated regime. We expect that the combination of currently available techniques and methodology from the fields of ion-atom interactions, in particular bichromatic dipole traps and optimized rf-based hybrid apparatus as well as protocols for the generation of ground-state polar molecules, is feasible within the scope of the envisioned studies. We identify and discuss several promising research directions in different regimes of the molecular ensembles, trap configurations and geometries. Such investigations include reactive inter-species processes with several open channels for chemical reactions and the creation of polyatomic molecular ions. Based on simple hypotheses, we expect that despite the seemingly much higher complexity compared to ion-atom or neutral molecule systems, the additional degrees of freedom, foremost the rotational states of the molecules, may be exploited and controlled to an extent allowing for the realization of molecule-ion collisions in the ultracold regime, as well as sympathetic cooling. Furthermore, we found indications that the interplay between attractive and repulsive interactions induced by the presence of the ion in a dense polar medium may play an important role in the dynamics of the many-body system.
@@
These capabilities open the way to access the sought-after regime of quantum interactions between ions and neutrals holding great promise to advance several fields and applications including refined control of collisional properties in ensembles of polar molecules, quantum chemistry, quantum-many-body physics, as well as quantum simulations and technologies.